Given that 𝑇 = 3√ 32𝑐4βˆ’π‘ž/2 . Show that 𝑐 = 1/2 4√2𝑇2 + π‘ž Β Β 


1
Expert's answer
2021-11-09T15:16:32-0500

"T = \sqrt[3]{32c^4-\frac{q}{2}}\\
=T^3 = 32c^4-\frac{q}{2}\\
=T^3+\frac{q}{2}=32c^4\\
=\frac{1}{32}(T^3+\frac{q}{2})=c^4\\
=\frac{1}{16}\cdot\frac{1}{2}(2T^3+q)=c^4\\
\implies c = \frac{1}{2}\sqrt[4]{\frac{1}{2}(2T^3+q)}"